OpenStudy (anonymous):

csc theta = -7/4, with theta in quadrant three. Find cot theta.

OpenStudy (anonymous):



OpenStudy (anonymous):

7^2 -4^2= 49-16=33 Sqrt[33]

OpenStudy (anonymous):

-Sqrt[33]/4